Greer Golf Supplies is an on line store that sells two types of golf balls: practice balls and tournament balls. The golf balls are sold in plastic sleeves containing three golf balls. practice balls sell for $4 per sleeve; tournament balls sell for $12 per sleeve. Owner carl rider purchases the golf balls directly from the manufacturer and pays $1 per sleeve for the practice balls and $4 per sleeve for the tournament balls. fixed costs total $14000 per month and include carl's salary website hosting and accounting and legal fees. when preparing the sales forecast for the year, carl assumed he would sell twice as many sleeves of practice balls as tournament balls.

Calculate the ANNUAL breakeven point for Greer Golf supplies using the sales forecast. what are the annual breakeven SALE$$ for practice sleeves? ( YOU DON'T NEED THE DATA TO CALCULATE THE ANSWER TO THIS QUESTION))
Answer #1
Practice Sleeve Tournament Sleeve Total
Sales price per unit 4.00 12.00
(-) Variable cost per unit 1.00 4.00
Contribution margin per unit 3.00 8.00
(x) Sales mix in units 2 1 3
Contribution margin 6.00 8.00 14.00
Weighted-average contribution margin per unit ($14/3) 4.67

Annual breakeven point = Fixed costs/Weighted average contribution margin per unit = $(14000*12)/$4.67 = 36,000 sleeves

Breakeven number of units for Tournament sleeves = 1/3 x 2998 = 36,000 = 12,000 sleeves

Breakeven sales for Practice sleeves = 12,000 * 4 = 96,000